# Decoherence of Histories: Chaotic Versus Integrable Systems

### v2

@@ Abstract
We study the emergence of decoherent histories in isolated systems based on exact numerical integration of the Schrödinger equation for a Heisenberg chain. We reveal that the nature of the system, which we switch from (i) chaotic to (ii) interacting integrable to (iii) non-interacting integrable, strongly impacts decoherence. From a finite size scaling law we infer a strong exponential suppression of coherences for (i), a weak exponential suppression for (ii) and no exponential suppression for (iii) on a relevant short (nonequilibrium) time scale. Moreover, for longer times we find stronger decoherence for (i) but the opposite for (ii), hinting even at a possible power-law decay for (ii) at equilibrium time scales. This behaviour is encoded in the multi-time properties of the quantum histories and it can not be explained by environmentally induced decoherence. Our results suggest that chaoticity plays a crucial role in the emergence of classicality in finite size systems.
Introduction.—The decoherence functional (DF)—quantifying the (de)coherence between different paths or histories in an isolated quantum system (precisely defined below)—plays a crucial role to explain the emergence of classicality [1, 2, 3, 4]. Fundamentally, it provides the condition for the existence of records of past events [5, 6, 7, 8, 9, 10], which may be used to identify branches in the universal wave function [1, 11], and it implies macrorealism (Leggett-Garg inequalities) [12, 13]. Practically, the DF determines whether a complex quantum process can be simulated by a simpler classical stochastic process [14, 15, 16, 17, 18, 19, 20, 21].
@@
The main contribution of this letter is an approximation-free evaluation of the DF for a realistic many-body system (a Heisenberg chain) based on exact numerical integration of the Schrödinger equation [28] without using classical ensembles. By leveraging the power of modern computers we extract a finite size scaling law and reveal that the nature of the system (chaotic, interacting integrable or free) strongly influences the emergence of decoherence (at least for finite size systems). Our results support an old conjecture by van Kampen [29] and significantly extend a few related studies evaluating the DF for pure states and finite size systems numerically [30, 31, 32, 33, 34, 19, 20, 11, 35] or on a quantum computer [36].
@@
Figure 1: $\langle A\rangle(t)$ as a function of rescaled time for initial states $|\psi^{+}_{0}\rangle$ (solid line) and $|\psi^{-}_{0}\rangle$ (dashed line) for the chaotic (i), interacting integrable (ii) and non-interacting integrable (iii) cases. The dotted (dash-dotted) vertical line represents $T=T_{\text{neq}}$ ($T=T_{\text{eq}}$), where $T_{\text{neq}}$ indicates the time at which $\langle A\rangle(t)$ decays to $e^{-1}$ of its initial value and $T_{\text{eq}}=20T_{\text{neq}}$.
@@
where $s^{\ell}_{x,y,z}=\sigma^{\ell}_{x,y,z}/2$ are spin operators at lattice sites $\ell$, $N$ is the length of the chain, and we assume periodic boundary conditions. Crucial for our purposes is that Eq. (8) contains three classes of systems for different parameter regimes: (i) for $\Delta_{1}\neq 0\neq\Delta_{2}$ (we choose $\Delta_{1}=1.5$, $\Delta_{2}=0.5$) the model is non-integrable (or chaotic) meaning that nearest-level-spacing follows a Wigner-Dyson distribution [60], and it satisfies the eigenstate thermalization hypothesis [61, 62]; (ii) for $\Delta_{1}\neq 0$ but $\Delta_{2}=0$ (we choose $\Delta_{1}=1.5$) it is an interacting integrable model, which is solvable by Bethe ansatz [63]; (iii) for $\Delta_{1}=\Delta_{2}=0$ the model is non-interacting integrable (or free) and can be mapped to a quadratic Hamiltonian (a set of free fermions) via Jordan-Wigner and Bogoliubov transformation [64]. The fact that (ii) is integrable but can not be mapped to free fermions like (iii) is crucial: it qualitatively influences its transport behavior [65], operator complexity [66] and, as we reveal below, its decoherence.
@@
$\Pi_{+}=\sum_{a_{k}\geq 0}|a_{k}\rangle\langle a_{k}|\quad\text{and}\quad\Pi_{-}=\sum_{a_{k}<0}|a_{k}\rangle\langle a_{k}|.$ (10)
@@
Numerical results.—To get an overall picture of the average dynamics we plot the expectation value $\langle A\rangle(t)$ of the coarse spin imbalance as a function of time in Fig. 1. This is done for two different non-equilibrium initial states $|\psi_{0}^{\pm}\rangle$, where $|\psi_{0}^{\pm}\rangle$ is a Haar random state restricted to the subspace ${\cal{H}}_{\pm}$. Remarkably, the behaviour in case (i) and (ii) agrees quantitatively: the system relaxes exponentially to its thermal equilibrium value $\langle A\rangle_{\text{eq}}=0$ with an equilibration time scale $\propto N^{2}$. In contrast, in case (iii) $\langle A\rangle(t)$ decays on a time scale $\propto N$ and fluctuates around $\langle A\rangle_{\text{eq}}$ without any clearly visible equilibration (up to the time that we considered).
@@
Decoherence is investigated in Figs. 2 and 3 for Haar random initial states $|\psi_{0}\rangle$. We plot in double logarithmic scale $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus the Hilbert space dimension $D$ for histories of lengths $L\in\{3,4,5\}$; the case $L=2$ has a universal typical decay owing to the Haar random nature of the initial state [19] as exemplified in the supplemental material (SM). The plots are obtained for constant time intervals $t_{k}-t_{k-1}=T$ for two different $T$: a nonequilibrium time scale $T_{\text{neq}}$ in Fig. 2 (identical to the dotted line in Fig. 1) and an equilibrium time scale $T_{\text{eq}}$ in Fig. 3 (identical to the dash-dotted line in Fig. 1). More precisely, $T_{\text{neq}}$ is defined as the time at which $\langle A\rangle(t)$ decays to $e^{-1}$ of its initial value [67] and the equilibrium time scale is defined as $T_{\text{eq}}=20T_{\text{neq}}$. While in the free model (iii) there is no clearly visible equilibration in Fig. 1, we use the same convention for $T_{\text{eq}}$ for better comparison. Finally, each data point in Figs. 2 and 3 is obtained by averaging $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over $2^{30-N}$ different realizations of $|\psi_{0}\rangle$. In the SM we show that the variance of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ decays as $D^{-1}$ for all three cases (i), (ii) and (iii) (likely as a consequence of dynamical typicality [68, 69]), which justifies our focus on the average behaviour.
@@
We then observe in Figs. 2 and 3 the following (an explanation follows later). First, for the chaotic case (i) we find a scaling law of the form $D^{-\alpha}$ (note that $D$ scales exponentially with the number of spins $N$) with $\alpha\approx 0.35$ ($\alpha\approx 0.25$) for $\overline{\epsilon}$ ($\epsilon_{\text{max}}$) at the nonequilibrium time scale and with $\alpha\approx 0.5$ for both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ at the equilibrium time scale. This indicates a robust exponential suppression (with respective to system size $N$) of coherences in chaotic systems. For the interacting integrable case (ii) we find three major differences compared to (i). First, the exponent $\alpha$ is notably smaller in all cases. Second, for $T_{\text{eq}}$ $\alpha$ is smaller than for $T_{\text{neq}}$, conversely to case (i). Indeed, for $T_{\text{eq}}$ (Fig. 3) we even find $\alpha\approx 0$, indicating a possible sub-exponential or power-law suppression (with respect to $N$) of coherences. Third, the exponent $\alpha$ for $\epsilon_{\text{max}}$ is roughly half the magnitude than $\alpha$ for $\overline{\epsilon}$ at $T_{\text{neq}}$, indicating stronger fluctuations in the DF among different pairs $(\boldsymbol{x};\boldsymbol{y})$ of histories. Finally, the free case (iii) is characterized for $T_{\text{neq}}$ by much larger coherences and for $T_{\text{eq}}$ by strong fluctuations among different $N$, making it unjustified to even speak of any scaling law $D^{-\alpha}$.
@@
Figure 4: Decoherence as a function of the time step $T=t_{n}-t_{n-1}$ for history length $L=5$. (a,b) $\overline{\epsilon},\epsilon_{\text{max}}$ for $N=24$ on a double logarithmic scale. (c,d) Scaling exponent $\alpha$ for $\overline{\epsilon},\epsilon_{\text{max}}$ obtained for system sizes $N=18,20,\ldots,28$ and as a mean over $2^{30-N}$ realizations of a Haar random $|\psi_{0}\rangle$. Red dashed line in (c,d) indicates $\alpha=0.5$.
@@
A particularly intriguing observation is the distinctive behaviour of decoherence as a function of $T$, which is further studied in Fig. S2. We first observe that decoherence is forbidden for $T\rightarrow 0$ owing to the quantum Zeno effect. Still, decoherence sets in on a short nonequilibrium time scale for (i) and (ii) and later for (iii). However, what we find surprising is that decoherence becomes weaker again with increasing $T$ for (ii) and (iii) but not for (i). Moreover, the scaling exponent $\alpha$ clearly shows a non-trivial behaviour as a function of $T$.
@@
Explanation.—We note that there is some consensus about the qualitative origin of decoherence (whether in the histories or EID framework), namely that coarse and slow (or “quasi-conserved”) observables of many-body systems decohere. However, other qualitative questions (e.g., is non-integrability essential or not?) have not been addressed in the past, and useful estimates of the DF are hard to obtain [19, 20]. Nevertheless, at least some quantitative features of the DF for chaotic systems seem to have a transparent explanantion.
@@
To this end, recall that the overlap $\langle\phi|\chi\rangle$ between two Haar random vectors $|\phi\rangle$ and $|\chi\rangle$ scales like $1/\sqrt{D}$ and that the subspace dimensions $V_{x}$ are proportional to $D$ for many relevant coarse-grainings. Thus, for equilibrium time scales $T_{\text{eq}}$ the history states $|\psi(\boldsymbol{x})\rangle$ in a chaotic system behave like randomly drawn typical states because they had time to explore the available Hilbert space in an unbiased way owing to the absence of conserved quantities (besides energy). Instead, for nonequilibrium time scales $T_{\text{neq}}$ the history states $|\psi(\boldsymbol{x})\rangle$ will not look fully randomized as they contain further information compared to the equilibrium case. Therefore, one expects the exponent $\alpha=\alpha(T)$ to obey $\alpha(T)<0.5$ for $T<T_{\text{eq}}$, but Fig. S2 reveals exceptions for short time windows. This indicates that the precise time dependence of $\alpha(T)$ is determined by $H$, $\{\Pi_{x}\}$ and $|\psi_{0}\rangle$ in a complicated way.
@@
Unfortunatey, it is even more complicated to explain the behaviour for cases (ii) and (iii). Certainly, owing to the extensive number of conserved quantities, the states $|\psi(\boldsymbol{x})\rangle$ can not explore the available Hilbert space in an unbiased fashion, which causes deviations from the behaviour of typical states. Yet, why the exponent $\alpha(T)$ becomes even smaller for larger $T$ remains a mystery. A speculative guess could be that at short time scales the dynamics is governed by time-dependent perturbation theory (Fermi’s golden rule), which is more determined by the rough structure of the spectrum of the Hamiltonian. At long times, instead, the fine structure of the Hamiltonian (its conserved quantities) become increasingly important for the dynamical behaviour.
@@
Conclusion.—We numerically extracted finite size scaling laws for the DF of a realistic quantum many-body system in an approximation-free way and we revealed decisive differences depending on the nature of the system. The chaotic case (i) showed a strong and robust emergence of decoherence in contrast to the interacting integrable case (ii) with a quantitatively much weaker form of decoherence. Note that this difference could not have been guessed from the quantitatively almost identical single-time behaviour shown in Fig. 1. A qualitative even weaker form of decoherence was observed for free models (iii), making it even hard to speak of any definite signature of decoherence. Clearly, our finite size calculations do not directly invalidate conclusions obtained from free models in the thermodynamic limit [22, 23, 6, 7, 25, 26, 27]. In particular, for another model (energy exchanges in an Ising chain, studied in the SM) we found a clear signature of decoherence also for the free case (while still being much weaker compared to the chaotic case).
@@
$\epsilon(\boldsymbol{x};\boldsymbol{y})=\frac{r_{\boldsymbol{x},\boldsymbol{y}}}{D^{\alpha}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for }\boldsymbol{x}\neq\boldsymbol{y}.$ (11)
@@
The notable difference in the value $\alpha$ at equilibrium time scales (and its time dependence) between chaotic and integrable systems suggests that $\alpha$ can be used as an indicator of quantum chaos. A main advantage of it is the applicability to system sizes beyond the reach of exact diagonalization, e.g., through real-time propagation methods such as the Chebyshev polynomial algorithm used here.
@@
Figure S2: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus Hilbert space dimension $D$ for history length $L=5$ for chaotic (a,d), interacting integrable (b,e) and non-interacting integrable (c,f) cases for time steps $T=T_{\text{neq}},2T_{\text{neq}},3T_{\text{neq}},T_{\text{eq}}$ (from light to dark blue). The dashed line indicate the scaling $D^{-1/2}$. The system size is $N=18,20,\dots,28$.
In addition to Fig. 4 in the main text, we further study the distinctive behaviour of decoherence as a function of $T$ in Fig. S2. Whereas in case (i) $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ consistently decrease and the exponent grows to $\alpha\approx 0.5$ for increasing $T$, the opposite happens for case (ii): decoherence becomes consistently weaker for larger $T$ and no exponential scaling is visible anymore for $\epsilon_{\text{max}}$ at $T_{\text{eq}}$. Finally, for case (iii) no clear pattern is visible and the large fluctuations make it difficult to speak of any scaling at all.
Next, we consider the variance of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ for $T_{\text{neq}}$ in Fig. S3 and $T_{\text{eq}}$ in Fig. S4. This is extracted from a histogram of the $2^{30-N}$ different realizations of $|\psi_{0}\rangle$. As we can see in all three cases (i), (ii) and (iii), the variance is orders of magnitude smaller than the mean and scales roughly like $1/D$, likely as a consequence of dynamical typicality. This is the reason why we can focus on the averaged values in the main text.
Figure S3: Variance of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus Hilbert space dimension $D$ for cases (i) (dark blue disks), (ii) (medium blue diamonds) and (iii) (light blue triangles) for $L=3,4,5$. The dashed line indicates the scaling $\propto D^{-1}$. The time step is $T=T_{\text{neq}}$ and the system sizes are $N=18,20,\dots,28$.
@@
Furthermore, we directly compare in Fig. S5 how the average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ coherence changes as a function of $L$ for the chaotic case. We observe an almost constant behaviour for $\overline{\epsilon}$ if $T=T_{\text{eq}}$ and only some mild changes for $T_{\text{neq}}$, which could be a result of the finite amount of samples. For $\epsilon_{\text{max}}$ instead we consistenly observe (slightly) larger values for larger $L$ for both $T_{\text{eq}}$ and $T_{\text{neq}}$. This is most likely caused by the fact that the DF has more entries for larger $L$ (the number of non-trivial entries grows like $M^{2L-1}-M^{L}$). Since $\epsilon_{\text{max}}$ was defined to measure statistical outliers, it is clear that there are bigger chances for stronger deviations if there are more elements to sample from.
In addition to the averaged and maximum value of $\epsilon_{\boldsymbol{x},\boldsymbol{y}}$, we also study its distribution. As an example, in Fig. S6, we show the distribution of the real and imaginary part of $\epsilon_{\boldsymbol{x},\boldsymbol{y}}$ for $N=20,\ L=5,\ T=T_{\text{eq}}$, where we exclude pairs of $\boldsymbol{x},\boldsymbol{y}$ for which $x_{n}\neq y_{n}$ or $x_{0}\neq y_{0}$. Data from $2^{10}$ different initial random states are taken into account. In the chaotic case (i), a similar shape of distribution is found for the real ($f_{R}(\epsilon)$) and imaginary part ($f_{I}(\epsilon)$). However, slight deviations are clearly observed, especially in the variance (second moments), indicating that $\epsilon_{\boldsymbol{x},\boldsymbol{y}}$ can not be random numbers in a strict sense. This is due to the fact that in a real system, starting from the same initial state, different histories are nevertheless correlated. However, in comparison with the integrable cases (ii) and (iii), where we observe large deviations between $f_{R}(\epsilon)$ and $f_{I}(\epsilon)$, the correlations in the chaotic case are almost negligible.

### v3

@@ Abstract
We study the emergence of decoherent histories in isolated systems based on exact numerical integration of the Schrödinger equation for a Heisenberg chain. We reveal that the nature of the system, which we switch from (i) chaotic to (ii) interacting integrable to (iii) non-interacting integrable, strongly impacts decoherence of coarse spin observables. From a finite size scaling law we infer a strong exponential suppression of coherences for (i), a weak exponential suppression for (ii) and no exponential suppression for (iii) on a relevant short (nonequilibrium) time scale. Moreover, for longer times we find stronger decoherence for (i) but the opposite for (ii), hinting even at a possible power-law decay for (ii) at equilibrium time scales. This behaviour is encoded in the multi-time properties of the quantum histories and it can not be explained by environmentally induced decoherence. Our results suggest that chaoticity plays a crucial role in the emergence of classicality in finite size systems.
Introduction.—Decoherent histories play a crucial role to explain the emergence of classicality in isolated quantum systems [1, 2, 3, 4]. Mathematically, the diagonality of the decoherence functional (DF)—quantifying the (de)coherence between different paths or histories in an isolated quantum system (precisely defined below)— provides the condition for the existence of records of past events [5, 6, 7, 8, 9, 10], which may be used to identify branches in the universal wave function [1, 11]. It also implies the validity of Leggett-Garg inequalities [12, 13] and allows to replace a complex quantum process by a simpler classical stochastic process [14, 15, 16, 17, 18, 19, 20, 21].
@@
The main contribution of this letter is an approximation-free evaluation of the DF for a realistic many-body system (a Heisenberg chain) based on exact numerical integration of the Schrödinger equation [28] without using classical ensembles, thereby significantly extending a few previous studies evaluating the DF for pure states and finite size systems numerically [29, 30, 31, 32, 33, 19, 20, 11, 34] or on a quantum computer [35]. By leveraging the power of modern computers we extract a finite size scaling law and reveal that the nature of the system (chaotic, interacting integrable or free) strongly influences the emergence of decoherence at least for finite size systems. Our results support van Kampen’s old idea that coarse and slow observables of chaotic quantum systems are described by classical stochastic processes [36]. Interestingly, while it has been argued that coarse and slow (or “quasi-conserved”) observables are important for decoherent histories [6, 37, 38, 39], the influence of chaos has never been considered.
@@
Figure 1: $\langle A\rangle(t)$ as a function of rescaled time for initial states $|\psi^{+}_{0}\rangle$ (solid line) and $|\psi^{-}_{0}\rangle$ (dashed line) for the chaotic (i), interacting integrable (ii) and non-interacting integrable (iii) cases for increasing system size (light to dark blue). The dotted (dash-dotted) vertical line represents $T=T_{\text{neq}}$ ($T=T_{\text{eq}}$), where $T_{\text{neq}}$ indicates the time at which $\langle A\rangle(t)$ decays to $e^{-1}$ of its initial value and $T_{\text{eq}}=20T_{\text{neq}}$.
@@
where $s^{\ell}_{x,y,z}=\sigma^{\ell}_{x,y,z}/2$ are spin operators at lattice sites $\ell$, $N$ is the length of the chain (chosen to be even), and we assume periodic boundary conditions. Crucial for our purposes is that Eq. (8) contains three classes of systems for different parameter regimes: (i) for $\Delta_{1}\neq 0\neq\Delta_{2}$ (we choose $\Delta_{1}=1.5$, $\Delta_{2}=0.5$) the model is non-integrable (or chaotic) meaning that nearest-level-spacing follows a Wigner-Dyson distribution [63], and it satisfies the eigenstate thermalization hypothesis [64, 65]; (ii) for $\Delta_{1}\neq 0$ but $\Delta_{2}=0$ (we choose $\Delta_{1}=1.5$) it is an interacting integrable model, which is solvable by Bethe ansatz [66]; (iii) for $\Delta_{1}=\Delta_{2}=0$ the model is non-interacting integrable (or free) and can be mapped to a quadratic Hamiltonian (a set of free fermions) via Jordan-Wigner transformation [67]. The fact that (ii) is integrable but can not be mapped to free fermions like (iii) is crucial: it qualitatively influences its transport behavior [68], operator complexity [69] and, as we reveal below, its decoherence.
@@
$\Pi_{+}=\sum_{a_{k}>0}|a_{k}\rangle\langle a_{k}|\quad\text{and}\quad\Pi_{-}=\sum_{a_{k}\leq 0}|a_{k}\rangle\langle a_{k}|.$ (10)
@@
Numerical results.—To get an overall picture of the average dynamics we plot the expectation value $\langle A\rangle(t)$ of the coarse spin imbalance as a function of time in Fig. 1 for two non-equilibrium initial states $|\psi_{0}^{\pm}\rangle$, where $|\psi_{0}^{\pm}\rangle$ is a Haar random state restricted to the subspace ${\cal{H}}_{\pm}$. Remarkably, the behaviour in case (i) and (ii) is very similar: the system relaxes approximately exponentially to its thermal equilibrium value $\langle A\rangle_{\text{eq}}=0$ with an equilibration time scale $\propto N^{2}$. In contrast, in case (iii) $\langle A\rangle(t)$ decays on a time scale $\propto N$ and fluctuates around $\langle A\rangle_{\text{eq}}$ without any clearly visible equilibration (up to the time that we considered).
Decoherence is investigated in Figs. 2 and 3 for Haar random initial states $|\psi_{0}\rangle$  [^1] We also considered initial product states and found results qualitatively similar to Figs. 2 and 3 (not shown here for brevity).. We plot in double logarithmic scale $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus the Hilbert space dimension $D$ for histories of lengths $L\in\{3,4,5\}$; the case $L=2$ has a universal typical decay owing to the Haar random nature of the initial state as exemplified in the supplemental material (SM). The plots are obtained for constant time intervals $t_{k}-t_{k-1}=T$ for two different $T$: a nonequilibrium time scale $T_{\text{neq}}$ in Fig. 2 (identical to the dotted line in Fig. 1) and an equilibrium time scale $T_{\text{eq}}$ in Fig. 3 (identical to the dash-dotted line in Fig. 1). More precisely, $T_{\text{neq}}$ is defined as the time at which $\langle A\rangle(t)$ decays to $e^{-1}$ of its initial value [71] and the equilibrium time scale is defined as $T_{\text{eq}}=20T_{\text{neq}}$. While in the free model (iii) there is no clearly visible equilibration in Fig. 1, we use the same convention for $T_{\text{eq}}$ for better comparison. Each data point in Figs. 2 and 3 is obtained by averaging $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over $2^{32-N}$ different realizations of $|\psi_{0}\rangle$. In the SM we also study the fluctuations of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ as a function of $|\psi_{0}\rangle$, which do not influence our conclusions below. In particular, the standard deviation scales like $D^{-1/2}$, likely as a consequence of dynamical typicality [72, 73].
@@
We then observe in Figs. 2 and 3 the following (an explanation follows later). First, for the chaotic case (i) we find a scaling law of the form $D^{-\alpha}$ (note that $D$ scales exponentially with the number of spins $N$) with $\alpha\approx 0.35$ ($\alpha\approx 0.25$) for $\overline{\epsilon}$ ($\epsilon_{\text{max}}$) at the nonequilibrium time scale and with $\alpha\approx 0.5$ for both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ at the equilibrium time scale. This indicates a robust exponential suppression (with respect to system size $N$) of coherences in chaotic systems. For the interacting integrable case (ii) we find three major differences compared to (i). First, the exponent $\alpha$ is notably smaller in all cases. Second, for $T_{\text{eq}}$, $\alpha$ is smaller than for $T_{\text{neq}}$, conversely to case (i). Indeed, for $T_{\text{eq}}$ (Fig. 3) we even find $\alpha\approx 0$, indicating a possible sub-exponential or power-law suppression (with respect to $N$) of coherences. Third, the exponent $\alpha$ for $\epsilon_{\text{max}}$ is roughly half the magnitude than $\alpha$ for $\overline{\epsilon}$ at $T_{\text{neq}}$, indicating stronger fluctuations in the DF among different pairs $(\boldsymbol{x};\boldsymbol{y})$ of histories. Finally, the free case (iii) is characterized for $T_{\text{neq}}$ by much larger coherences and for $T_{\text{eq}}$ by strong fluctuations among different $N$, making it unjustified to even speak of any scaling law $D^{-\alpha}$.
@@
Figure 4: Decoherence as a function of the time step $T=t_{k}-t_{k-1}$ for history length $L=5$. (a,b) $\overline{\epsilon},\epsilon_{\text{max}}$ for $N=26$ on a double logarithmic scale. (c,d) Scaling exponent $\alpha$ for $\overline{\epsilon},\epsilon_{\text{max}}$ obtained for system sizes $N=18,20,\ldots,28$ and as a mean over $2^{31-N}$ realizations of a Haar random $|\psi_{0}\rangle$. Black dotted (dash-dotted) line indicates $T_{\text{neq}}$ $(T_{\text{eq}})$. Red dashed line in (c,d) indicates $\alpha=0.5$.
@@
A particularly intriguing observation is the distinctive behaviour of decoherence as a function of $T$, which is studied in detail in Fig. 4. We first observe that decoherence is forbidden for $T\rightarrow 0$ owing to the quantum Zeno effect. Still, decoherence sets in on a short nonequilibrium time scale for (i) and (ii) and later for (iii). What we found surprising is that for longer times $T\geq T_{\text{neq}}$ decoherence becomes weaker with increasing $T$ for (ii) and (iii) but not for (i). The scaling exponent $\alpha$ saturates to $\alpha\approx 0.5$ for (i), it decays to approximately zero for (ii), and it exhibits strong fluctuations for (iii).
@@
Explanation.—We note that there is some consensus about the qualitative origin of decoherence (whether in the histories or EID framework), namely that coarse and slow observables of many-body systems decohere, as considered here (in addition, we study in detail the connection to quasi-conserved hydrodynamic modes of spin density wave operators in the SM). However, other qualitative questions (e.g., is non-integrability essential or not?) have not been addressed in the past, and useful estimates of the DF are hard to obtain [19, 20]. Nevertheless, at least some quantitative features of the DF for chaotic systems seem to have a transparent explanation.
@@
To this end, recall that the overlap $\langle\phi|\chi\rangle$ between two Haar random vectors $|\phi\rangle$ and $|\chi\rangle$ scales like $1/\sqrt{D}$ and that the subspace dimensions $V_{x}$ are proportional to $D$ for many relevant coarse-grainings. Thus, for equilibrium time scales $T_{\text{eq}}$ the history states $|\psi(\boldsymbol{x})\rangle$ in a chaotic system behave like randomly drawn typical states because they had time to explore the available Hilbert space in an unbiased way owing to the absence of conserved quantities (besides energy), and the irregularity of the time-dependent phases makes recoherences extremely unlikely for the vast majority of times. Instead, for nonequilibrium time scales $T_{\text{neq}}$ the history states $|\psi(\boldsymbol{x})\rangle$ will not look fully randomized as they contain further information compared to the equilibrium case. Therefore, one expects the exponent $\alpha=\alpha(T)$ to obey $\alpha(T)<0.5$ for $T<T_{\text{eq}}$, but Fig. 4 reveals exceptions for short time windows, which we can not explain at the moment. This indicates that the precise time dependence of $\alpha(T)$ is determined by $H$, $\{\Pi_{x}\}$ and $|\psi_{0}\rangle$ in a complicated way.
@@
Unfortunately, it is even more complicated to explain the behaviour for cases (ii) and (iii). Certainly, owing to the extensive number of conserved quantities, the states $|\psi(\boldsymbol{x})\rangle$ can not explore the available Hilbert space in an unbiased fashion, which causes deviations from the behaviour of typical states. Yet, that the exponent $\alpha(T)$ becomes even smaller for larger $T$ appears counterintuitive. We believe the reason is that the dynamics, which is determined by terms of the form $\exp(-i\omega_{jk}T)$ with $\omega_{jk}$ a spectral gap, is less sensitive to the fine structure of $\{\omega_{jk}\}$ for short times $T$. The interesting point is that in the borderline case (ii) this sensitivity is revealed by the DF (a higher order correlation functions) but not by $\langle A\rangle(t)$ (cf. Fig. 1).
@@
Conclusion.—We numerically extracted finite size scaling laws for the DF of a realistic quantum many-body system in an approximation-free way and we revealed decisive differences depending on the nature of the system. The chaotic case (i) showed a strong and robust emergence of decoherence in contrast to the interacting integrable case (ii) with a quantitatively much weaker form of decoherence. Note that this difference could not have been guessed from the quantitatively almost identical single-time behaviour shown in Fig. 1. A qualitative even weaker form of decoherence was observed for free models (iii), making it even hard to speak of any definite signature of decoherence. Clearly, our finite size and model dependent calculations do not directly invalidate conclusions obtained from free models in the thermodynamic limit [22, 23, 6, 7, 25, 26, 27]. However, for another model (energy exchanges in an Ising chain, studied in the SM) we also found strong differences between the chaotic and free case, suggesting that great care is needed when assessing the decoherence of free models.
@@
$\epsilon(\boldsymbol{x};\boldsymbol{y})=\delta_{\boldsymbol{x},\boldsymbol{y}}+(1-\delta_{\boldsymbol{x},\boldsymbol{y}})\frac{r_{\boldsymbol{x},\boldsymbol{y}}}{D^{\alpha}}.$ (11)
@@
Finally, the notable difference in the value $\alpha$ at equilibrium time scales (and its time dependence) between chaotic and integrable systems suggests that $\alpha$ can be used as an indicator of quantum chaos. A main advantage of it is the applicability to system sizes beyond the reach of exact diagonalization, e.g., through real-time propagation methods such as the Chebyshev polynomial algorithm used here.
@@
We here report on further numerical results that strengthen our main message. We start by supplementing additional results for the XXZ Heisenberg chain studied in the main text (subsection S1), continued by comparing the chaotic and free case for energy imbalance in an Ising chain (subsection S2), and finally we look at the level statistics to demonstrate that our models are chaotic (subsection S3).
@@
Figure S2: Inverse SNR versus Hilbert space dimension $D$ for cases (i) (dark blue disks), (ii) (medium blue diamonds) and (iii) (light blue triangles) for $L=3,4,5$. The time step is $T=T_{\text{neq}}$ and the system sizes are $N=18,20,\dots,28$.
@@
Next, we consider the question whether it is justified to focus on the averages of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over the $2^{32-N}$ different realizations of $|\psi_{0}\rangle$ (denoted here as $m_{\overline{\epsilon}}$ and $m_{\epsilon_{\text{max}}}$), as done in the main text. To this end, we computed the standard deviation $v_{\overline{\epsilon}}$ and $v_{\epsilon_{\text{max}}}$, which we found to scale like $D^{-1/2}$ in all cases (i), (ii) and (iii) and for all $L\in\{3,4,5\}$ (likely as a consequence of dynamical typicality). This is not shown here for brevity because it is more revealing to consider the inverse signal-to-noise ratio (SNR) $v_{\overline{\epsilon}}/m_{\overline{\epsilon}}\sim D^{\alpha(\overline{\epsilon})-1/2}$ and $v_{\epsilon_{\text{max}}}/m_{\epsilon_{\text{max}}}\sim D^{\alpha(\epsilon_{\text{max}})-1/2}$ for $T_{\text{neq}}$ in Fig. S2 and for $T_{\text{eq}}$ in Fig. S3. Since the integrable dynamics (ii) and (iii) is characterized by very weak decoherence (if at all), their inverse SNR quickly becomes vanishingly small. This is not so for the chaotic dynamics (i), where at equilibrium time scales $\alpha\approx 1/2$ implies no vanishing of the inverse SNR with system size. However, we can clearly see in Fig. S3 that the standard deviation is about an order of magnitude smaller than the mean value, justifying our focus on the average in the main text. Moreover, we also find that $v_{\overline{\epsilon}}/m_{\overline{\epsilon}}$ shows a clear decrease with $L$ because the size of the DF increases with $L$.
Furthermore, we directly compare in Fig. S4 how the average $\overline{\epsilon}$ and maximum coherence $\epsilon_{\text{max}}$ changes as a function of $L$ for the chaotic case. We observe in all four cases rather mild changes as a function of $L$, and there seems to be a convergence to a common value in unison with the scaling law extracted in the main text (however, for $L\gg 1$, which was numerically not accessible for us, “recoherences” must eventually happen). Moreover, for $\epsilon_{\text{max}}$ we consistently observe a slight increase for larger $L$ for both $T_{\text{eq}}$ and $T_{\text{neq}}$. Since $\epsilon_{\text{max}}$ was defined to measure statistical outliers, this is likely caused by the fact that the DF has more entries for larger $L$ (the number of non-trivial entries grows like $M^{2L-1}-M^{L}$).
In addition to the averaged and maximum value of $\epsilon_{\boldsymbol{x},\boldsymbol{y}}$, we also study its distribution. As an example, in Fig. S5, we show the distribution of the real and imaginary part of $\epsilon_{\boldsymbol{x},\boldsymbol{y}}$ for $N=20,\ L=5,\ T=T_{\text{eq}}$, where we exclude pairs of $\boldsymbol{x},\boldsymbol{y}$ for which $x_{n}\neq y_{n}$ or $x_{0}\neq y_{0}$. Data from $2^{12}$ different initial Haar random states are taken into account. In the chaotic case (i), a similar shape of distribution is found for the real ($f_{R}(\epsilon)$) and imaginary part ($f_{I}(\epsilon)$). However, slight deviations are clearly observed, especially in the variance (second moments), indicating that $\epsilon_{\boldsymbol{x},\boldsymbol{y}}$ can not be random numbers in a strict sense. This is due to the fact that in a real system, starting from the same initial state, different histories are nevertheless correlated. However, in comparison with the integrable cases (ii) and (iii), where we observe large deviations between $f_{R}(\epsilon)$ and $f_{I}(\epsilon)$, the correlations in the chaotic case are almost negligible.